# Enhancing Depression-Diagnosis-Oriented Chat with Psychological State Tracking

### v1

@@ 4.3 Fine-tuning
In this work, we fine-tune ChatGLM3 111https://github.com/THUDM/ChatGLM3 with a parameter efficient approach, i.e., Low-Rank Adaptation (LoRA). LoRA maintains the weights of pre-trained LMs while introducing trainable rank decomposition matrices into each transformer layer, making it feasible to fine-tune LLMs with much fewer computational resources.
We fine-tune an LLM to track the psychological state and generate the response generation jointly, given the crafted example and the annotated label. Specifically, the objective is to predict the next token based on language modeling:
@@ 5 Experiments
Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 | Topic ACC. |
Transformer- | 7.28% | 0.21 | 0.1570 | 0.29 | - |
ChatGPT- | 10.01% | 0.19 | 0.3254 | 0.15 | - |
BART- | 19.29% | 0.35 | 0.2866 | 0.09 | - |
CPT- | 19.79% | 0.36 | 0.2969 | 0.07 | - |
Ours- | 30.85% | 0.47 | 0.4524 | 0.17 | - |
Transformer | 13.43% | 0.33 | 0.2620 | 0.04 | 36.82% |
ChatGPT | 17.52% | 0.30 | 0.4080 | 0.19 | 14.62% |
BART | 28.62% | 0.48 | 0.4053 | 0.07 | 59.56% |
CPT | 29.40% | 0.48 | 0.4142 | 0.06 | 59.77% |
Ours | 39.76% | 0.50 | 0.5305 | 0.11 | 56.98% |
Transformer* | 25.37% | 0.41 | 0.3905 | 0.04 | - |
ChatGPT* | 29.46% | 0.42 | 0.5158 | 0.23 | - |
BART* | 37.02% | 0.54 | 0.4920 | 0.07 | - |
CPT* | 37.45% | 0.54 | 0.4943 | 0.06 | - |
Ours* | 45.28% | 0.56 | 0.5794 | 0.10 | - |
5.1 Baselines
We leverage the original Transformer architecture Vaswani et al. (2017) fine-tuned on MedDialog dataset Zeng et al. (2020) as a foundational model. BART model Lewis et al. (2019) and the CPT model Shao et al. (2021) are employed as our primary baselines, with the same configuration as previous studies Yao et al. (2022). ChatGPT (gpt-3.5-turbo) 222https://openai.com/blog/chatgpt is listed as a baseline as well, which is optimized for chat by introducing Reinforcement Learning with Human Feedback (RLHF).
5.2 Implementation Details
Our model is implemented in PyTorch, using ChatGLM3-6b as the base architecture. In the supervised fine-tuning process, we apply LoRA to all linear layers of the model, where LoRA rank is set to 64. We set the batch size, max context length, and learning rate to 32, 1024, and 2e-4 respectively. The model is trained on one A800 GPU for 5 epochs, which costs about 8 hours.
@@ 5.3 Automatic Evaluation
Response Generation
Metrics like BLEU-2 Papineni et al. (2002), Rouge-L Lin (2004) and METEOR Banerjee and Lavie (2005) are employed to assess the response generation quality. In addition, we calculate DIST-2 Li et al. (2015) to demonstrate the diversity of responses. Tokenization is performed using the jieba333https://github.com/fxsjy/jieba library, following the same approach as in D.
Table 2 shows the results of the automatic evaluation for the response generation task. We have the following observations. First, without incorporating POSTs, the model achieves considerably better performances on BLEU-2, ROUGE-L and METEOR. It is mainly attributed to the larger scale of the language model and optimization for Chinese dialogue generation. Second, the integration of POSTs markedly enhanced diagnostic response capabilities, particularly in BLEU-2 and METEOR, by effectively assessing the diagnosis stage, summarizing patient information, and inferring potential diagnoses. By tracking psychological states, the POST guides the model in generating more rational and interpretable responses, and a further enhancement is observed while the golden POSTs are given. Third, The results from ChatGPT indicate relatively lower BLEU-2 and ROUGE-L scores due to the lack of dataset-specific fine-tuning. The introduction of POSTs notably improved ChatGPT’s performance by 7%, with substantial improvements upon using golden POSTs, underscoring the effectiveness of the POST strategy.
Summarization
Table 3 presents the results for dialog summary task. We provide the model with all the dialogue context and concatenate it with the patient’s portrait information. The model then generates a summary of the patient’s symptoms. The results show that our model surpasses both BART and CPT in text generation and also exhibits better diversity according to the DIST-2 scores.
Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 |
BART | 16.44% | 0.26 | 0.25 | 0.19 |
CPT | 16.45% | 0.26 | 0.24 | 0.21 |
Ours | 18.32% | 0.30 | 0.29 | 0.26 |
Severity Classification
Results for depression severity and suicide risk severity classification are listed in Tables 4 and 5. To be consistent with D, we use average weighted precision, recall, and F1 scores computed by sklearn444https://scikit-learn.org to evaluate classification results.
The results demonstrate that our approach outperforms the baseline models in most scenarios, with a notable performance enhancement in 4-classification tasks when utilizing summaries compared to dialog history. However, the difference is less pronounced in the 2-classification tasks due to the increased complexity of the 4-classification tasks. Summaries provide concise and effective diagnostic information about the patient’s condition, which leads to a more substantial improvement in performance. Additionally, the variance in the 4-classification results based on dialog history is relatively high, which might be partly due to uneven label distribution in the dataset. Besides, the 2-classification accuracy of suicide severity based on dialog history have reached 90%, indicating this approach’s practical viability for diagnosing suicidal tendencies from dialog history.
Task | Input | Model | Precision | Recall | F1 |
2-class | dialog | BERT | 0.81 | 0.80 | 0.80 |
BART | 0.80 | 0.79 | 0.79 | ||
CPT | 0.79 | 0.78 | 0.78 | ||
Ours | 0.88 | 0.86 | 0.84 | ||
summary | BERT | 0.90 | 0.90 | 0.90 | |
BART | 0.89 | 0.89 | 0.89 | ||
CPT | 0.92 | 0.92 | 0.92 | ||
Ours | 0.91 | 0.91 | 0.91 | ||
4-class | dialog | BERT | 0.49 | 0.45 | 0.45 |
BART | 0.53 | 0.53 | 0.52 | ||
CPT | 0.49 | 0.47 | 0.46 | ||
Ours | 0.57 | 0.56 | 0.53 | ||
summary | BERT | 0.67 | 0.66 | 0.66 | |
BART | 0.68 | 0.67 | 0.66 | ||
CPT | 0.73 | 0.72 | 0.72 | ||
Ours | 0.79 | 0.76 | 0.76 |
Task | Input | Model | Precision | Recall | F1 |
2-class | dialog | BERT | 0.81 | 0.78 | 0.79 |
BART | 0.77 | 0.75 | 0.75 | ||
CPT | 0.84 | 0.82 | 0.82 | ||
Ours | 0.91 | 0.90 | 0.91 | ||
4-class | dialog | BERT | 0.72 | 0.64 | 0.66 |
BART | 0.70 | 0.66 | 0.65 | ||
CPT | 0.76 | 0.68 | 0.70 | ||
Ours | 0.79 | 0.71 | 0.73 |
@@ 5.4 Human Evaluation
As shown in Table 6, our model surpasses the baseline in the majority of aspects, particularly in doctor-likeness and engagingness, indicating significant improvements compared to other models. This suggests that incorporating the POST enhances our model’s ability to emulate doctors’ interactions, capturing the patient’s state in real-time, and allowing for flexible topic transitions throughout the dialogue. However, we also note that our model did not outperform ChatGPT in terms of comforting. This could be attributed to ChatGPT’s ability to generate extensive comforting and advisory responses without specific tuning. The dialogue examples of these models can be found in Appendix B.
Comparisons | Aspect | Win | Lose | Tie |
Ours vs. CPT | Flu. | 44.7 | 21.2 | 34.1 |
Com. | 68.9 | 15.2 | 15.9 | |
Doc. | 62.1 | 14.4 | 23.5 | |
Eng. | 56.8 | 11.4 | 31.8 | |
Ours vs. ChatGPT | Flu. | 39.4 | 34.8 | 25.8 |
Com. | 33.3 | 55.3 | 11.4 | |
Doc. | 53.8 | 34.1 | 12.1 | |
Eng. | 41.7 | 28.8 | 29.5 |
@@ 5.5 Ablation Studies
Model | BLEU-2 | ROUGE-L | METEOR | DIST-2 |
Ours* | 45.28% | 0.56 | 0.5794 | 0.10 |
@@
The study reveals that omitting the next topic (w/o Next) significantly impacts response generation, as it is directly correlated with the subsequent response’s relevance and effectiveness. This omission led to an increase in the DIST-2 value, suggesting a reduction in response content diversity due to the focused nature of topic-specific responses. Furthermore, excluding the stage (w/o Stage) notably affects the quality of generation, indicating the importance of the diagnostic dialogue phase in tailoring responses to be stage-appropriate and rational. In contrast, the absence of information (w/o Info) and summary (w/o Sum) components showed lesser impact on response generation quality, due to their roles in summarizing dialogue history rather than directly influencing responses. Nonetheless, these elements are crucial for their interpretability and utility in clinical settings, aiding doctors in summarizing patient symptoms and providing preliminary diagnostic insights, thereby facilitating further diagnosis and having a substantial role in clinical psychological consultations.
@@ 6 Conclusions
In this paper, we incorporate Psychological State Tracking (POST) within LLM to guide response generation for doctors during depression diagnosis consultations. In particular, the state is defined based on the Albert Ellis’ ABC Model in psychology, which illuminates a profound connection between patients’ information changes and doctors’ strategic planning. Extensive experiments show that the integration of psychological state tracking significantly enhances the performance of LLMs to generate responses in depression-diagnosis-oriented chat. Furthermore, our approach also provides explicit interpretations for using appropriate strategies in different situations to collect information or comfort patients for depression diagnosis. In the future, we will explore more specialized and fine-grained state tracking methods and incorporate patient personalized information to guide diagnosis-oriented chat.
Limitations
The main limitation of our approach lies in the evaluation metrics. In the field of diagnostic chat, traditional automatic evaluation methods may not accurately reflect the true effectiveness of the models. As highlighted by Liu et al. (2016), there exists a certain discrepancy between the results and human evaluations. Developing more comprehensive, considerate, and task-specific evaluation metrics remains a challenge to be addressed in future research.
Ethics Statement
In this study, we annotate the psychological state collaboratively by professional labors and GPT-4, followed by a thorough data cleaning and filtering process. This step ensured the removal of any private or sensitive information from the dataset.
Acknowledgements
We acknowledge the contributions of the SJTU X-LANCE Lab555https://x-lance.sjtu.edu.cn/en in constructing the original depression-diagnosis-oriented chat dataset (i.e., D). We cite the work and conduct further research by adapting the dataset and exploring LLM-based method from the psychological views.
Appendix A Prompt Template
Description: I want you to serve as a psychological counselor, and use rational emotive behavior therapy and ABC theory to mark the current patient situation and the next strategy. Please mark the input dialogues according to the requirements and examples in Chinese. The sequence numbers of the input dialogues must strictly correspond one-to-one with those of the outputs, with no merging or omissions allowed. Requirements: Definitions and marking requirements of four parts of POST: <Stage> The current diagnostic stage (Activating Event (A) / Belief (B) / Consequence (C), please enter the complete stage, and answers like "same as above" are not allowed). <Info> Current patient information (keep information from historical conversations as briefly as possible). <Sum> Diagnosis result of the patient’s condition (you need to refine it to the underlying disease). <Next> Next step strategy (choose the most suitable option from 11 strategies). Examples: Examples pre-labeled by humans Input: Input data Output:
Appendix B Case Study
Figures 5, 6, 7 respectively display dialogue examples of our model, CPT, and ChatGPT under the same patient background, where these models play the role of doctors. In the diagram, statements of empathetic consolation are highlighted with green wavy lines, while inquiries of diagnosis are underlined in blue. It is evident from the illustration that, in comparison to other models, our model offers more natural empathetic consolation and topics of inquiry that are closely aligned with the diagnostic state. The responses from CPT tend to be relatively rigid, with less empathetic consolation and more straightforward, singular diagnostic inquiries. Although ChatGPT provides commendable empathetic consolation, its replies often are too lengthy, containing extensive analysis and advice, yet they do not closely correlate with the diagnostic content, lacking the distinctive characteristics of a doctor.

### v2

@@ 4.3 Fine-tuning
In this work, we fine-tune large language models with a parameter-efficient approach, i.e., Low-Rank Adaptation (LoRA). LoRA maintains the weights of pre-trained LMs while introducing trainable rank decomposition matrices into each transformer layer, making it feasible to fine-tune LLMs with much fewer computational resources.
We fine-tune an LLM to track the psychological state and generate the response jointly, given the crafted example and the annotated label. Specifically, the objective is to predict the next token based on language modeling:
@@ 5 Experiments
5.1 Experimental Setups
5.1.1 Baselines
We leverage the CPT model [19] as our primary baseline, as it achieved the best performance in previous studies [25]. And we use the same configuration as them. ChatGPT (gpt-3.5-turbo) 222https://openai.com/blog/chatgpt is also listed as a baseline, which performs well in most tasks.
5.1.2 Implementation Details
Our model uses ChatGLM3-6b333https://github.com/THUDM/ChatGLM3 as the base architecture, and is implemented in PyTorch. During the supervised fine-tuning process, we apply LoRA to all linear layers of the model, where LoRA rank is set to 64. We set the batch size, max context length, and learning rate to 32, 1024, and 2e-4, respectively. The model is trained on one A800 GPU for 5 epochs, which costs about 8 hours.
[…]
ChatGPT | / | 10.01% | 0.19 | 0.3254 | 0.15 | - |
CPT | / | 19.79% | 0.36 | 0.2969 | 0.07 | - |
ChatGLM3 | / | 30.85% | 0.47 | 0.4524 | 0.17 | - |
ChatGPT | +POST | 17.52% | 0.30 | 0.4080 | 0.19 | 14.62% |
+7.51% | +0.11 | +0.0826 | +0.04 | - | ||
CPT | +POST | 34.15% | 0.44 | 0.4609 | 0.06 | 57.12% |
+14.36% | +0.08 | +0.1640 | -0.01 | - | ||
ChatGLM3 | +POST | 39.76% | 0.50 | 0.5305 | 0.11 | 56.98% |
+8.91% | +0.03 | +0.0781 | -0.06 | - | ||
ChatGPT | +POST* | 29.46% | 0.42 | 0.5158 | 0.23 | - |
+19.45% | +0.23 | +0.1904 | +0.08 | - | ||
CPT | +POST* | 41.94% | 0.55 | 0.5285 | 0.04 | - |
+22.15% | +0.19 | +0.2316 | -0.03 | - | ||
ChatGLM3 | +POST* | 45.28% | 0.56 | 0.5794 | 0.10 | - |
+14.43% | +0.09 | +0.1270 | -0.07 | - | ||
[…]
@@ 5.2 Automatic Evaluation
We evaluate the performance of turn-based response generation given the dialogue history in the dataset. Typical automatic metrics for text generation like BLEU-2 [16], Rouge-L [12] and METEOR [1] are employed to assess the response generation quality. In addition, we calculate DIST-2 [11] to demonstrate the diversity of generated responses.
Table 2 shows the results of the automatic evaluation for the response generation task. We have the following observations. First, all models show substantial improvements in BLEU-2, ROUGE-L, and METEOR after applying POST, indicating that this method can significantly enhance the generation quality. Second, the application of POST resulted in considerable performance boosts across all models, with CPT showing the most significant improvement of +14.36% in BLEU-2 over its baseline. This enhancement is attributed to the framework’s ability to effectively assess the diagnostic stage, summarize patient information, and infer potential diagnoses. Further improvements are observed when golden POSTs are given. Third, as ChatGPT cannot be fine-tuned for our specific task, its baseline performance in this domain is restricted. However, the introduction of POST notably improved ChatGPT’s performance by +7.51% in BLEU-2, with substantial improvements upon using golden POST (+19.45%), demonstrating the effectiveness of the POST strategy. Interestingly, while ChatGPT shows improved DIST-2 with POST (+0.04), both CPT and ChatGLM3 exhibit slight decreases in DIST-2 (-0.01 and -0.06 respectively). This divergence reflects that unguided baselines naturally produce more diverse outputs, while POST’s structure constrains variety. ChatGPT’s unique characteristics allow it to maintain diversity even with structured guidance.
@@ 5.3 Human Evaluation
[…]
ChatGLM3 +POST vs. ChatGLM3 | Flu. | 63.4 | 32.5 | 4.1 |
Com. | 95.9 | 4.1 | 0.0 | |
Doc. | 92.7 | 6.5 | 0.8 | |
Eng. | 93.5 | 5.7 | 0.8 | |
[…]
As shown in Table 3, the baseline using the POST method performs better in most aspects, with particularly significant improvements in Comforting, Doctor-likeness, and Engagingness. Regarding fluency, the gap between our two models is not as large as other metrics, since the method has a relatively minor impact on fluency. These results suggest that incorporating the POST enhances our model’s ability to emulate doctors’ interactions, capturing the patient’s state in real-time, and allowing for flexible strategy transitions throughout the dialogue.
@@ 5.4 Ablation Studies
The study reveals that omitting the next strategy (w/o Next) significantly impacts response generation, as it is directly correlated with the subsequent response’s relevance and effectiveness. Furthermore, excluding the stage (w/o Stage) notably affects the quality of generation, indicating the importance of the diagnostic dialogue phase in tailoring responses to be stage-appropriate and rational. In contrast, the absence of information (w/o Info) and summary (w/o Sum) components showed less impact on response generation quality, due to their roles in summarizing dialogue history rather than directly influencing responses. Nonetheless, these elements are crucial for their interpretability and utility in clinical settings, aiding doctors in summarizing patient symptoms and providing preliminary diagnostic insights, thereby facilitating further diagnosis and having a substantial role in clinical psychological consultations.
[…]
ChatGLM3 +POST* | 45.28% | 0.56 | 0.5794 | 0.10 |
@@
[…]
@@ 6 Conclusions
In this paper, we incorporate Psychological State Tracking (POST) within LLM to guide response generation for doctors during depression diagnosis consultations. In particular, the state is defined based on Albert Ellis’ ABC Model in psychology, which illuminates a profound connection between patients’ information changes and doctors’ strategic planning. Extensive experiments show that the integration of psychological state tracking significantly enhances the performance of LLMs to generate responses in depression-diagnosis-oriented chat. Furthermore, our approach also provides explicit interpretations for using appropriate strategies in different situations to collect information or comfort patients for depression diagnosis. In the future, we will explore more specialized and fine-grained state tracking methods and incorporate patient personalized information to guide diagnosis-oriented chat.
6.0.1 Acknowledgement
This research is funded by the National Nature Science Foundation of China (No. 62477010 and No.62307028), the Natural Science Foundation of Shanghai (No. 23ZR1441800), Shanghai Science and Technology Innovation Action Plan (No. 24YF2710100 and No.23YF1426100) and Shanghai Special Project to Promote High-quality Industrial Development (No. 2024-GZL-RGZN-02008).